# Sample-Specific Noise Injection For Diffusion-Based Adversarial Purification

### v1

@@ 1 Introduction
Deep neural networks (DNNs) are vulnerable to adversarial examples, which is a longstanding problem in deep learning (Szegedy et al., 2014; Goodfellow et al., 2015). Adversarial examples aim to mislead DNNs into making erroneous predictions by adding imperceptible adversarial noise to clean examples, which pose a significant security threat in critical applications (Dong et al., 2019; Finlayson et al., 2019; Cao et al., 2021; Jing et al., 2021). To defend against adversarial examples, adversarial purification (AP) stands out as a representative defensive mechanism, by leveraging pre-trained generative models to purify adversarial examples back towards their natural counterparts before feeding into a pre-trained classifier (Yoon et al., 2021; Nie et al., 2022). Notably, AP methods benefit from their modularity, as the purifier operates independently of the downstream classifier, which facilitates seamless integration into existing systems and positions AP as a practical approach to improve the adversarial robustness of DNN-based classifiers.
Recently, diffusion-based purification (DBP) methods have gained much attention as a promising framework in AP, which leverage the denoising nature of diffusion models to mitigate adversarial noise (Nie et al., 2022; Xiao et al., 2023; Lee & Kim, 2023). Generally, diffusion models train a forward process that maps from data distributions to simple distributions, e.g., Gaussian, and reverse this mapping via a reverse generative process (Ho et al., 2020; Song et al., 2021b). When applied in DBP methods, the forward process gradually injects Gaussian noise into the input sample, while the reverse process gradually purify noisy sample to recover the clean sample. The quality of the purified sample heavily depends on the amount of Gaussian noise added to the input during the forward process, which can be controlled by a noise level parameter $t^{*}$. Existing DBP methods (Nie et al., 2022; Xiao et al., 2023; Lee & Kim, 2023) manually select a constant $t^{*}$ for all samples.
@@
To implement SSNI, inspired by the fact that scores (i.e., $\nabla_{{\mathbf{x}}}{\log p_{t}({\mathbf{x}})}$) reflect the directional momentum of samples toward the high-density areas of clean data distribution (Song & Ermon, 2019), we use score norms (i.e., $\left\|\nabla_{{\mathbf{x}}}{\log p_{t}({\mathbf{x}})}\right\|$) as a natural metric to measure the deviation of a data point from the clean data distribution. In Section 3, we establish the relationship between the score norm and the noise level required for different samples. Specifically, samples with different score norms tend to have accumulated different noise levels. Furthermore, we empirically show that the cleaner samples – those closer to the clean data distribution – exhibit lower score norm, justifying the rationale of using score norms for reweighting $t^{*}$. Concretely, we use a pre-trained score network to estimate the score norm for each sample. Based on this, we propose two reweighting functions that adaptively adjust $t^{*}$ according to its score norm, achieving sample-specific noise injections (see Section 4.3). Notably, this reweighting process is lightweight, ensuring that SSNI is computationally feasible and can be applied in practice with minimal overhead (see Section 5.6).
Through extensive evaluations on benchmark image datasets such as CIFAR-10 (Krizhevsky et al., 2009) and ImageNet-1K (Deng et al., 2009), we demonstrate the effectiveness of SSNI in Section 5. Specifically, combined with different DBP methods (Nie et al., 2022; Xiao et al., 2023; Lee & Kim, 2023), SSNI can boost clean accuracy and robust accuracy simultaneously by a notable margin against the well-designed adaptive white-box attack (see Section 5.2).
@@ 2 Preliminary and Related Work
Score and score norm. In diffusion models, score refers to the gradient of the log-probability density $\nabla_{{\mathbf{x}}}{\log p_{t}({\mathbf{x}})}$, which maps the steepest ascent direction of the log-density in the probability space (Song & Ermon, 2019). The associated score norm $\left\|\nabla_{{\mathbf{x}}}{\log p({\mathbf{x}})}\right\|$ measures the gradient’s magnitude, reflecting how much a data point deviates from the clean data distribution: points located in low-probability regions typically exhibit larger score norms, while those situated closer to high-density regions of clean data manifest smaller score norms. This property is valuable for adversarial example detection, as demonstrated by Yoon et al. (2021) who establish that score norms can effectively distinguish between adversarial and clean examples. However, direct computation of the score is often intractable for complex data distributions (e.g., images). In practice, it is estimated using neural networks $s_{\theta}:{\mathcal{X}}\times{\mathbb{R}}_{+}\to{\mathcal{X}}$ that take as input both a data point ${\mathbf{x}}$ and a specified timestep $t^{\rm S}\in{\mathbb{R}}_{+}$ to approximate $\nabla_{{\mathbf{x}}}{\log p_{t}({\mathbf{x}})}$ by minimizing the Fisher divergence between the true and estimated scores. Also, $s_{\theta}({\mathbf{x}},t^{\rm S})$ enables efficient computation of $\left\|\nabla_{{\mathbf{x}}}{\log p({\mathbf{x}})}\right\|$.
Diffusion-based purification. Adversarial purification (AP) leverages generative models as an add-on module to purify adversarial examples before classification. Within this context, diffusion-based purification (DBP) methods have emerged as a promising framework, exploiting the inherent denoising nature of diffusion models to filter out adversarial noise (Nie et al., 2022; Wang et al., 2022; Xiao et al., 2023; Lee & Kim, 2023). Specifically, DBP works with pre-trained diffusion model by first corrupting an adversarial input through the forward process and then iteratively reconstructing it via the reverse process. This projects the input back onto the clean data manifold while stripping away adversarial artifacts. Wang et al. (2022) introduce input guidance during the reverse diffusion process to ensure the purified outputs stay close to the inputs. Lee & Kim (2023) propose a fine-tuned gradual noise scheduling for multi-step purifications. Bai et al. (2024) improve the reverse diffusion process by incorporating contrastive objectives.
@@ 3 Motivation
Score norms vary across perturbation budgets. Adversarial and clean examples are from distinct distributions (Gao et al., 2021). Motivated by this, we further investigate how different perturbation budgets $\epsilon$ affect score norms under adversarial attacks (Figure 2). Specifically, we compute the score norm of different samples undergoing PGD+EOT $\ell_{\infty}(\epsilon=8/255)$ with perturbation budgets varying between 00 and $8/255$ on CIFAR-10. We observe a consistent pattern: score norms scale directly with perturbation strength: larger $\epsilon$ values lead to higher norms, whereas smaller perturbations (i.e., cleaner samples) lead to lower norms The findings extend the role of score norms from differentiating adversarial/clean samples (Yoon et al., 2021) to differentiate adversarial examples based on their perturbation strength.
@@ 4.1 Framework of SSNI
$h({\mathbf{x}})=C\circ R\circ D({\mathbf{x}}),\quad\text{with }{\mathbf{x}}={\mathbf{x}}_{0}.$ (3)
@@ 4.3 Realization of SSNI
where $p_{t}({\mathbf{x}})$ is the marginal probability density and $t^{\rm S}$ is the maximum noise level for EPS. EPS computes the expectation of the scores of perturbed images across different noise levels $t\sim U(0,t^{\rm S})$, making it more robust to the changes in noise levels. Notably, this $t^{\rm S}$ is different from the sample-shared noise injection level $t^{*}$. We will omit $t^{\rm S}$ from the notation of ${\rm EPS}({\mathbf{x}})$ for brevity hereafter. Following Zhang et al. (2023), we set $t^{\rm S}=20$. In practice, a score $\nabla_{\mathbf{x}}\log_{p_{t}}({\mathbf{x}})$ can be approximated by a score network. Specifically, we employ a score network pre-trained using the score matching objective (Song & Ermon, 2019).
@@
WRN-28-10 Nie et al. (2022) 89.71$\pm$0.72 47.98$\pm$0.64
+ SSNI-N 93.29$\pm$0.37 (+3.58) 48.63$\pm$0.56 (+0.65)
Wang et al. (2022) 92.45$\pm$0.64 36.72$\pm$1.05
+ SSNI-N 94.08$\pm$0.33 (+1.63) 40.95$\pm$0.65 (+4.23)
Lee & Kim (2023) 90.10$\pm$0.18 56.05$\pm$1.11
+ SSNI-N 93.55$\pm$0.55 (+3.45) 56.45$\pm$0.28 (+0.40)
WRN-70-16 Nie et al. (2022) 90.89$\pm$1.13 52.15$\pm$0.30
+ SSNI-N 94.47$\pm$0.51 (+3.58) 52.47$\pm$0.66 (+0.32)
Wang et al. (2022) 93.10$\pm$0.51 43.55$\pm$0.58
+ SSNI-N 95.57$\pm$0.24 (+2.47) 46.03$\pm$1.33 (+2.48)
Lee & Kim (2023) 89.39$\pm$1.12 56.97$\pm$0.33
+ SSNI-N 93.82$\pm$0.24 (+4.43) 57.03$\pm$0.28 (+0.06)
@@
WRN-28-10 Nie et al. (2022) 91.80$\pm$0.84 82.81$\pm$0.97
+ SSNI-N 93.95$\pm$0.70 (+2.15) 82.75$\pm$1.01 (-0.06)
Wang et al. (2022) 92.45$\pm$0.64 82.29$\pm$0.82
+ SSNI-N 94.08$\pm$0.33 (+1.63) 82.49$\pm$0.75 (+0.20)
Lee & Kim (2023) 90.10$\pm$0.18 83.66$\pm$0.46
+ SSNI-N 93.55$\pm$0.55 (+3.45) 84.05$\pm$0.33 (+0.39)
WRN-70-16 Nie et al. (2022) 92.90$\pm$0.40 82.94$\pm$1.13
+ SSNI-N 95.12$\pm$0.58 (+2.22) 84.38$\pm$0.58 (+1.44)
Wang et al. (2022) 93.10$\pm$0.51 85.03$\pm$0.49
+ SSNI-N 95.57$\pm$0.24 (+2.47) 84.64$\pm$0.51 (-0.39)
Lee & Kim (2023) 89.39$\pm$1.12 84.51$\pm$0.37
+ SSNI-N 93.82$\pm$0.24 (+4.43) 84.83$\pm$0.33 (+0.32)
@@ 5.1 Experimental Settings
Datasets and model architectures. We consider two datasets for our evaluations: CIFAR-10 (Krizhevsky et al., 2009), and ImageNet-1K (Deng et al., 2009). For classifiers, we use the pre-trained WideResNet-28-10 and WideResNet-70-16 for CIFAR-10, and the pre-trained ResNet-50 for ImageNet-1K. For diffusion models, we employ two off-the-shelf diffusion models trained on CIFAR-10 and ImageNet-1K (Song et al., 2021b; Dhariwal & Nichol, 2021).
@@
Baseline settings. We use three well-known DBP methods as our baselines: DiffPure (Nie et al., 2022), GDMP (Wang et al., 2022) and GNS (Lee & Kim, 2023). The detailed configurations can be found in Appendix C. For the reverse process within diffusion models, we consider DDPM sampling method (Ho et al., 2020) in the DBP methods.
Evaluation settings for DBP baselines. Following Lee & Kim (2023), we use a fixed subset of 512 randomly sampled images for all evaluations due to high computational cost of applying adaptive white-box attacks to DBP methods. Lee & Kim (2023) provide a robust evaluation framework for existing DBP methods and demonstrate that PGD+EOT (Madry et al., 2018; Athalye et al., 2018b) is the golden standard for DBP evaluations. Therefore, following Lee & Kim (2023), we mainly use adaptive white-box PGD+EOT attack with 200 PGD iterations for CIFAR-10 and 20 PGD iterations for ImageNet-1K. We use 20 EOT iterations for all experiments to mitigate the stochasticity introduced by the diffusion models. As PGD is a gradient-based attack, we compute the gradients of the entire process from a surrogate process. The details of the surrogate process is explained in Appendix D. We also evaluate DBP methods under adaptive BPDA+EOT attack (Athalye et al., 2018a), which leverages an identity function to approximate the direct gradient rather than direct computing the gradient of the defense system.
Evaluation settings for SSNI. Since SSNI introduces an extra reweighting process than DBP baselines, we implicitly design two adaptive white-box attacks by considering the entire defense mechanism of SSNI (i.e., adaptive white-box PGD+EOT attack and adaptive white-box BPDA+EOT attack). To make a fair comparison, we evaluate SSNI on adaptive white-box attacks with the same configurations mentioned above. The algorithmic descriptions for the adaptive white-box PGD+EOT attack and adaptive white-box BPDA+EOT attack is provided in Appendix E and F. In addition, to evaluate the generalization and adaptability of SSNI to diverse adversarial attacks, we further include AutoAttack (Croce & Hein, 2020), DiffAttack (Chen et al., 2024b) and the Diff-PGD attack (Xue et al., 2023) in Section 5.4. We set the iteration number to 5 for Diff-PGD.
@@ 5.4 Defending Against Additional Attacks
5.5 Ablation Study
Ablation study on $\tau$ in SSNI-N. We investigate how the temperature coefficient $\tau$ in Eq. (8) affects the performance of SSNI-N against adaptive white-box PGD+EOT $\ell_{\infty}(\epsilon=8/255)$ attack on CIFAR-10 in Figure 4. The temperature coefficient $\tau$ controls the sharpness of the curve of the non-linear reweighting function. A higher $\tau$ leads to a more smooth transition between the low and high values of the reweighting function, resulting in less sensitivity to the changes of the input. From Figure 4, the standard accuracy remains stable across different $\tau$s, while the robust accuracy increases to the climax when $\tau=20$. Therefore, we choose $\tau^{*}=20$ for the non-linear reweighting function to optimize the accuracy-robustness trade-off for DBP methods.
Ablation study on sampling methods. DiffPure originally used an adjoint method to efficiently compute the gradients of the system, but Lee & Kim (2023) and Chen et al. (2024a) suggest to replace adjoint solver with sdeint solver for the purpose of computing full gradients more accurately (Li et al., 2020; Kidger et al., 2021). Therefore, we investigate whether using different sampling methods affect the performance of DBP methods (here we use DiffPure as the baseline method). We further compare the results with DDIM sampling method (Song et al., 2021a), which is a faster sampling method than DDPM (Ho et al., 2020). From Table 5, DDPM achieves the best accuracy-robustness trade-off among the three sampling methods, and thus we select DDPM as the sampling method for all baseline methods.
@@ 5.6 Compute Resource
The inference time (in seconds) for incorporating SSNI modules into existing DBP methods on CIFAR-10 and ImageNet-1K are reported in Tables 6. The inference time is measured as the time it takes for a single test image to complete the purification process. Specifically, SSNI is approximately 0.5 seconds slower than baseline methods on CIFAR-10 and 5 seconds slower than baseline methods on ImageNet-1K. Thus, compared with DBP baseline methods, this reweighting process is lightweight, ensuring that SSNI is computationally feasible and can be applied in practice with minimal overhead. We implemented our code on Python version 3.8, CUDA version 12.2.0, and PyTorch version 2.0.1 with Slurm Workload Manager. We conduct each of the experiments on up to 4 $\times$ NVIDIA A100 GPUs.

### v2

@@ 1 Introduction
Deep neural networks (DNNs) are vulnerable to adversarial examples, which is a longstanding problem in deep learning (Szegedy et al., 2014; Goodfellow et al., 2015). Adversarial examples aim to mislead DNNs into making erroneous predictions by adding imperceptible adversarial noise to clean examples, which pose a significant security threat in critical applications (Dong et al., 2019; Cao et al., 2021; Jing et al., 2021; Han et al., 2025). To defend against adversarial examples, adversarial purification (AP) stands out as a representative defensive mechanism, by leveraging pre-trained generative models to purify adversarial examples back towards their natural counterparts before feeding into a pre-trained classifier (Yoon et al., 2021; Nie et al., 2022). Notably, AP methods benefit from their modularity, as the purifier operates independently of the downstream classifier, which facilitates seamless integration into existing systems and positions AP as a practical approach to improve the adversarial robustness of DNN-based classifiers.
Recently, diffusion-based purification (DBP) methods have gained much attention as a promising framework in AP, which leverage the denoising nature of diffusion models to mitigate adversarial noise (Nie et al., 2022; Xiao et al., 2023; Lee and Kim, 2023). Generally, diffusion models train a forward process that maps from data distributions to simple distributions, e.g., Gaussian, and reverse this mapping via a reverse generative process (Ho et al., 2020; Song et al., 2021b). When applied in DBP methods, the forward process gradually injects Gaussian noise into the input sample, while the reverse process gradually purify noisy sample to recover the clean sample. The quality of the purified sample heavily depends on the amount of Gaussian noise added to the input during the forward process, which can be controlled by a noise level parameter $t^{*}$. Existing DBP methods (Nie et al., 2022; Xiao et al., 2023; Lee and Kim, 2023) manually select a constant $t^{*}$ for all samples.
@@
To implement SSNI, inspired by the fact that scores (i.e., $\nabla_{{\mathbf{x}}}{\log p_{t}({\mathbf{x}})}$) reflect the directional momentum of samples toward the high-density areas of clean data distribution (Song and Ermon, 2019), we use score norms (i.e., $\left\|\nabla_{{\mathbf{x}}}{\log p_{t}({\mathbf{x}})}\right\|$) as a natural metric to measure the deviation of a data point from the clean data distribution. In Section 3, we establish the relationship between the score norm and the noise level required for different samples. Specifically, samples with different score norms tend to have accumulated different noise levels. Furthermore, we empirically show that the cleaner samples – those closer to the clean data distribution – exhibit lower score norm, justifying the rationale of using score norms for reweighting $t^{*}$. Concretely, we use a pre-trained score network to estimate the score norm for each sample. Based on this, we propose two reweighting functions that adaptively adjust $t^{*}$ according to its score norm, achieving sample-specific noise injections (see Section 4.3). Notably, this reweighting process is lightweight, ensuring that SSNI is computationally feasible and can be applied in practice with minimal overhead (see Section 5.6).
Through extensive evaluations on benchmark image datasets such as CIFAR-10 (Krizhevsky et al., 2009) and ImageNet-1K (Deng et al., 2009), we demonstrate the effectiveness of SSNI in Section 5. Specifically, combined with different DBP methods (Nie et al., 2022; Xiao et al., 2023; Lee and Kim, 2023), SSNI can boost clean accuracy and robust accuracy simultaneously by a notable margin against the well-designed adaptive white-box attack (see Section 5.2).
@@ 2 Preliminary and Related Work
Score and score norm. In diffusion models, score refers to the gradient of the log-probability density $\nabla_{{\mathbf{x}}}{\log p_{t}({\mathbf{x}})}$, which maps the steepest ascent direction of the log-density in the probability space (Song and Ermon, 2019). The associated score norm $\left\|\nabla_{{\mathbf{x}}}{\log p({\mathbf{x}})}\right\|$ measures the gradient’s magnitude, reflecting how much a data point deviates from the clean data distribution: points located in low-probability regions typically exhibit larger score norms, while those situated closer to high-density regions of clean data manifest smaller score norms. This property is valuable for adversarial example detection, as demonstrated by Yoon et al. (2021) who establish that score norms can effectively distinguish between adversarial and clean examples. However, direct computation of the score is often intractable for complex data distributions (e.g., images). In practice, it is estimated using neural networks $s_{\theta}:{\mathcal{X}}\times{\mathbb{R}}_{+}\to{\mathcal{X}}$ that take as input both a data point $\mathbf{x}$} and a specified timestep $t^{\rm S}\in{\mathbb{R}}_{+}$ to approximate $\nabla_{{\mathbf{x}}}{\log p_{t}({\mathbf{x}})}$ by minimizing the Fisher divergence between the true and estimated scores. Also, $s_{\theta}({\mathbf{x}},t^{\rm S})$ enables efficient computation of $\left\|\nabla_{{\mathbf{x}}}{\log p({\mathbf{x}})}\right\|$.
Diffusion-based purification. Adversarial purification (AP) leverages generative models as an add-on module to purify adversarial examples before classification. Within this context, diffusion-based purification (DBP) methods have emerged as a promising framework, exploiting the inherent denoising nature of diffusion models to filter out adversarial noise (Nie et al., 2022; Wang et al., 2022; Xiao et al., 2023; Lee and Kim, 2023). Specifically, DBP works with pre-trained diffusion model by first corrupting an adversarial input through the forward process and then iteratively reconstructing it via the reverse process. This projects the input back onto the clean data manifold while stripping away adversarial artifacts. Wang et al. (2022) introduce input guidance during the reverse diffusion process to ensure the purified outputs stay close to the inputs. Lee and Kim (2023) propose a fine-tuned gradual noise scheduling for multi-step purifications. Bai et al. (2024) improve the reverse diffusion process by incorporating contrastive objectives.
@@ 3 Motivation
Score norms vary across perturbation budgets. Adversarial and clean examples are from distinct distributions (Gao et al., 2021). Motivated by this, we further investigate how different perturbation budgets $\epsilon$ affect score norms under adversarial attacks (Figure 2). Specifically, we compute the score norm of different samples undergoing PGD+EOT $\ell_{\infty}(\epsilon=8/255)$ with perturbation budgets varying between 0 and $8/255$ on CIFAR-10. We observe a consistent pattern: score norms scale directly with perturbation strength: larger $\epsilon$ values lead to higher norms, whereas smaller perturbations (i.e., cleaner samples) lead to lower norms The findings extend the role of score norms from differentiating adversarial/clean samples (Yoon et al., 2021) to differentiate adversarial examples based on their perturbation strength.
@@ 4.1 Framework of SSNI
[…]
@@ 4.3 Realization of SSNI
where $p_{t}({\mathbf{x}})$ is the marginal probability density and $t^{\rm S}$ is the maximum noise level for EPS. EPS computes the expectation of the scores of perturbed images across different noise levels $t\sim U(0,t^{\rm S})$, making it more robust to the changes in noise levels. Notably, this $t^{\rm S}$ is different from the sample-shared noise injection level $t^{*}$. We will omit $t^{\rm S}$ from the notation of ${\rm EPS}({\mathbf{x}})$ for brevity hereafter. Following Zhang et al. (2023), we set $t^{\rm S}=20$. In practice, a score $\nabla_{\mathbf{x}}\log_{p_{t}}({\mathbf{x}})$ can be approximated by a score network. Specifically, we employ a score network pre-trained using the score matching objective (Song and Ermon, 2019).
@@
[…]
@@
[…]
@@ 5.1 Experimental Settings
Datasets and model architectures. We consider two datasets for our evaluations: CIFAR-10 (Krizhevsky et al., 2009), and ImageNet-1K (Deng et al., 2009). For classifiers, we use the pre-trained WideResNet-28-10 and WideResNet-70-16 for CIFAR-10, and the pre-trained ResNet-50 for ImageNet-1K. For diffusion models, we employ two off-the-shelf diffusion models trained on CIFAR-10 and ImageNet-1K (Song et al., 2021b; Dhariwal and Nichol, 2021).
@@
Baseline settings. We use three well-known DBP methods as our baselines: DiffPure (Nie et al., 2022), GDMP (Wang et al., 2022) and GNS (Lee and Kim, 2023). The detailed configurations can be found in Appendix C. For the reverse process within diffusion models, we consider DDPM sampling method (Ho et al., 2020) in the DBP methods.
Evaluation settings for DBP baselines. Following Lee and Kim (2023), we use a fixed subset of 512 randomly sampled images for all evaluations due to high computational cost of applying adaptive white-box attacks to DBP methods. Lee and Kim (2023) provide a robust evaluation framework for existing DBP methods and demonstrate that PGD+EOT (Madry et al., 2018; Athalye et al., 2018b) is the golden standard for DBP evaluations. Therefore, following Lee and Kim (2023), we mainly use adaptive white-box PGD+EOT attack with 200 PGD iterations for CIFAR-10 and 20 PGD iterations for ImageNet-1K. We use 20 EOT iterations for all experiments to mitigate the stochasticity introduced by the diffusion models. As PGD is a gradient-based attack, we compute the gradients of the entire process from a surrogate process. The details of the surrogate process is explained in Appendix D. We also evaluate DBP methods under adaptive BPDA+EOT attack (Athalye et al., 2018a), which leverages an identity function to approximate the direct gradient rather than direct computing the gradient of the defense system.
Evaluation settings for SSNI. Since SSNI introduces an extra reweighting process than DBP baselines, we implicitly design two adaptive white-box attacks by considering the entire defense mechanism of SSNI (i.e., adaptive white-box PGD+EOT attack and adaptive white-box BPDA+EOT attack). To make a fair comparison, we evaluate SSNI on adaptive white-box attacks with the same configurations mentioned above. The algorithmic descriptions for the adaptive white-box PGD+EOT attack and adaptive white-box BPDA+EOT attack is provided in Appendix E and F. In addition, to evaluate the generalization and adaptability of SSNI to diverse adversarial attacks, we further include AutoAttack (Croce and Hein, 2020), DiffAttack (Chen et al., 2024b) and the Diff-PGD attack (Xue et al., 2023) in Section 5.4. We set the iteration number to 5 for Diff-PGD.
@@ 5.5 Ablation Study
Ablation study on sampling methods. DiffPure originally used an adjoint method to efficiently compute the gradients of the system, but Lee and Kim (2023) and Chen et al. (2024a) suggest to replace adjoint solver with sdeint solver for the purpose of computing full gradients more accurately (Li et al., 2020; Kidger et al., 2021). Therefore, we investigate whether using different sampling methods affect the performance of DBP methods (here we use DiffPure as the baseline method). We further compare the results with DDIM sampling method (Song et al., 2021a), which is a faster sampling method than DDPM (Ho et al., 2020). From Table 5, DDPM achieves the best accuracy-robustness trade-off among the three sampling methods, and thus we select DDPM as the sampling method for all baseline methods.
@@ 5.6 Compute Resource
The inference time (in seconds) for incorporating SSNI modules into existing DBP methods on CIFAR-10 and ImageNet-1K are reported in Tables 6. The inference time is measured as the time it takes for a single test image to complete the purification process. Specifically, SSNI is approximately 0.5 seconds slower than baseline methods on CIFAR-10 and 5 seconds slower than baseline methods on ImageNet-1K. Thus, compared with DBP baseline methods, this reweighting process is lightweight, ensuring that SSNI is computationally feasible and can be applied in practice with minimal overhead. We implemented our code on Python version 3.8, CUDA version 12.2.0, and PyTorch version 2.0.1 with Slurm Workload Manager. We conduct each of the experiments on up to 4 $\times$ NVIDIA A100 GPUs (see https://github.com/tmlr-group/SSNI).